OpenStudy (nannie):

PLEASE show mw how to solve this: 8/v=9/v-1/3

OpenStudy (anonymous):

1/3=9/v-8/v 1/3=1/v v=3

OpenStudy (anonymous):

it would be negative

OpenStudy (lgbasallote):

multiply both sides by v so that it would transfer to the numerator 8 = 9 + v/3..then multiply both sides by 3 24 = 27 + v v = -3

OpenStudy (nannie):

Thank you!